Can anyone help with ANY of these? Please!

1.Complete each exercise leaving imaginary answer in i form.
3 squareroots of -24

2.Complete each exercise leaving imaginary answer in i form.
(3 - i) + (1 - 2i)

3.Complete each exercise leaving imaginary answer in i form.
(7 + 3i) - (2 +3i)

4.Complete each exercise leaving imaginary answer in i form.
The squareroot of -25 + the squareroot of -16 + the squareroot of -9
= ____

5.Complete each exercise leaving imaginary answer in i form.
(5 -i)(5 +i)

6.Complete each exercise leaving imaginary answer in i form.
(-3 - 5i)(-3 + 5i)

7.Complete each exercise leaving imaginary answer in i form.
5-6i^11

8.Complete each exercise leaving imaginary answer in i form.
10(4-3i)

9.Complete each exercise leaving imaginary answer in i form.
i^12 + 1

10.Complete each exercise leaving imaginary answer in i form.
(5-2i^2)^2

11.Complete each exercise leaving imaginary answer in i form.
6 / 2i

12.Complete each exercise leaving imaginary answer in i form.

(2 + the square root of -3) (2 - the squareroot of -3)
___________________________ ___________________________
(2) (2)
* It doesn't come out right when i submit it but, each equation is over 2.

13.Complete each exercise leaving imaginary answer in i form.

4 + i
________
3 - 2i


Answer by stanbon(75887)

1.Complete each exercise leaving imaginary answer in i form.
3 squareroots of -24
= 3sqrt(-24)
=3*2i*sqrt(6)
= 6sqrt(6)i
====================

2.Complete each exercise leaving imaginary answer in i form.
(3 - i) + (1 - 2i)
= 4 - 3i
=======================
3.Complete each exercise leaving imaginary answer in i form.
(7 + 3i) - (2 +3i)
= 5
=============================
4.Complete each exercise leaving imaginary answer in i form.
The squareroot of -25 + the squareroot of -16 + the squareroot of -9
= sqrt(-25) + sqrt(-16) + sqrt(-9)
---
= 5i + 4i + 3i
---
= 12i
=================================

5.Complete each exercise leaving imaginary answer in i form.
(5 -i)(5 +i)
= 5^2 - i^2
= 25 + 1
= 26
====================================
6.Complete each exercise leaving imaginary answer in i form.
(-3 - 5i)(-3 + 5i)
= (-3)^2 - (5i)^2
= 9 + 25
= 34
=====================================
7.Complete each exercise leaving imaginary answer in i form.
5-6i^11
= 6 -6*(-i)
= 6 + 6i
=====================================
8.Complete each exercise leaving imaginary answer in i form.
10(4-3i)
= 40 - 30i
=====================================
9.Complete each exercise leaving imaginary answer in i form.
i^12 + 1
= 1 + 1
= 2
========================================
10.Complete each exercise leaving imaginary answer in i form.
(5-2i^2)^2
= 25 - 20i^2 + (2i)^2
= 25 + 20 + 4
= 49
========================================
11.Complete each exercise leaving imaginary answer in i form.
6 / 2i
= [6/2i][2i/2i]
= 12i/-4
= -3i
=========================================
12.Complete each exercise leaving imaginary answer in i form.

(2 + the square root of -3) (2 - the squareroot of -3)
= 4 - -3
= 7
========================================
___________________________ ___________________________
(2) (2)
* It doesn't come out right when i submit it but, each equation is over 2.
13.Complete each exercise leaving imaginary answer in i form.
(4 + i)/(3-2i)
---
= [(4+i)(3-2i)]/[(3+2i)/(3+2i)]
---
= [(4+i)(3-2i)]/[5]
= [12+2 +3i-8i]/5
= [14-5i]/5
=========================
Cheers,
Stan H.